# Novel Direct Alpha Spectroscopy Technique for 225Ac Radiopharmaceutical detection in Cancer Cells

### v1

@@ 2.5.3 Direct Alpha Spectroscopy of Cell Cultures Incubated with Labeled Activity
Three sets of sample cups were prepared as described in Section 2.5. The first set contained [225Ac]Ac-crown-TATE, the second set contained [225Ac]Ac-PSMA-617 as a negative control (assuming the AR42J cells do not have receptors for PSMA-617), and the third set was incubated with [225Ac]Ac3+. For each set, a corresponding series of reference sample cups was prepared with the same amount of labeled activity and following the same procedure, but without cells. Each sample cup was placed on the detector, and spectra were acquired for 900 seconds, achieving a statistical uncertainty of less than 1%. The experiment was repeated three times, and the standard deviation was applied to the results. Figure 3 illustrates a schematic view of the experimental setup.
@@
Figure 3: Schematic view of the experimental setup for direct alpha spectroscopy, showing the arrangement of AR42J cells with an approximate diameter of 13 µm. The cells are placed on a 2.5 µm Mylar foil, with a 100 µm air gap separating the cell sample and the Si PIN diode (300 µm thickness). The media layer above the cells is approximately 3 mm thick. The yellow arrows represent the trajectory of emitted alpha particles.
@@ 2.5.4 Geant4 simulation
The silicon photodiode detector used in the simulation had dimensions of 18 × 18 mm and a thickness of 300 µm. A cylindrical sample cup with a diameter of 13 mm and a length of 3 mm was separated by a 100 µm air gap from the detector surface. A Mylar foil with a thickness of 2.5 µm was placed at the bottom of the sample cups. The 225Ac was randomly distributed within the aqueous solution inside the sample cups. By varying the thickness of the water layers, we explored how interaction depth affects the energy spectrum detected by the silicon photodiode, providing valuable insights into the behavior of 225Ac under different conditions. Two sets of simulations were performed: in one, the 225Ac was distributed on the Mylar foil, and in the other, the 225Ac was distributed in a 13 µm water solution depth, representing the cell diameter. Figure 4 shows the Geant4 simulation geometry. Simulated spectra are included in Figure 11 and Figure 12 in comparison to experimental results.
Figure 4: Geant4 simulation showing the trajectory of alpha particles emitted from a 225Ac source. The alpha particles are distributed in 13 µm water solution representing the cell diameter, passing through the Mylar foil. The blue lines represent the simulated paths of alpha particles, demonstrating their interactions and energy deposition within the detector setup.
@@ 3.1 Alpha decay spectrum from 225Ac calibration source
Figure 5 shows the alpha spectrum for 225A obtained using the BAD detection system, revealing four distinct peaks for 225Ac (5830 keV), 221Fr (6340 keV), 217At (7067 keV), and 213Po (8376 keV). The detector has been calibrated using the highest energy flank of each peak, while the tail observed on the lower energy side is attributed to the energy loss of alpha particles as they pass through the Mylar foil and the air gap before reaching the detector. Calibration was performed with a 225Ac source of known activity, dried directly onto the Mylar foil in the sample cup geometry, ensuring that energy shifts caused by the detector assembly (Mylar foil and air gap) are already accounted for. The overall detector efficiency across the entire energy range was calculated to be 37%.
@@ 3.2 Cell Culture Distribution
Figure 6 shows the distribution of AR42J cells on the Mylar foil. The cells appear to be evenly distributed across the surface of the sample cup, with no visible clumping or aggregation. This even distribution is crucial for ensuring uniform exposure during alpha spectroscopy. Microspheres with a known diameter of 20.1 $\mu$m were used as a reference, and the cell diameter was determined based on these microspheres using a microscope image and ImageJ software. A Gaussian fit was applied to the distribution, resulting in a cell diameter of 13.16 $\pm$ 0.12 $\mu$m. 7(a) shows the distribution of cell diameters in comparison with the microspheres.
The size of the cells is important because it ensures that alpha particles can exit the cells and reach the detector. An SRIM (Stopping and Range of Ions in Matter) simulation was performed to calculate the range of alpha particles in a water-based material, approximating the composition of the cells. 7(b) shows the Bragg peak from the SRIM simulation, indicating that the range of alpha particles in water is approximately 50 $\mu$m. Given that the cell diameter is 13 $\mu$m, the simulation confirms that alpha particles can traverse the cells and exit into the detector. This is critical for measuring the uptake of alpha-emitting radiopharmaceuticals.
@@
Figure 7: (a): Normalized distribution of microspheres and cells based on their diameter, with an inset showing a microscopic image of cells with reference spheres. Microspheres (black) have a diameter of 20 $\mu$m, while cells (red) show a wider distribution centered around 13.16 $\pm$ 0.12 $\mu$m. (b): Bragg peak showing the energy deposition of alpha particles in water, calculated using SRIM. The range of alpha particles in water is approximately 50 $\mu$m, confirming that alpha particles can exit AR42J cells (diameter $\sim$ 13 $\mu$m) and be detected by the detector.
@@ 3.3 Alpha spectrum from cell sample incubated with [225Ac]Ac-crown-TATE
8(a) presents the alpha spectrum of [225Ac]Ac-crown-TATE from the cell sample, alongside the reference and cell supernatant samples, acquired using the BAD detection system. 8(b) shows the logarithmic scale of the same spectrum. Although the detector resolution does not clearly distinguish between the 225Ac and 221Fr peaks, the 217At and 213Po peaks are still recognizable. The number of counts was normalized to the initial activity measured by a calibrated HPGe detector, enabling a direct comparison between the three sample types. Three replicates of each sample were prepared, and the standard deviation is reflected in the plot. As can be seen in the cell sample spectrum, all peaks exhibit noticeable broadening, which is due to the energy loss of alpha particles as they traverse through the cells before reaching the detector. The spectrum shows significantly lower activity in the supernatant compared to the cell sample, indicating a clear uptake of activity by the cells. A small 213Po peak is observed in the supernatant 8(b), which is likely due to the release of 213Bi from the radiopharmaceutical complex into the media. This release might occur as a result of the high recoil energy generated during the alpha decay of 225Ac, which can disrupt the coordination bonds within the radiopharmaceutical complex. Such disruption may lead to the dissociation of progeny isotopes, such as 213Bi, from the original complex, allowing them to migrate into the surrounding media. The redistribution of daughter isotopes can affect both the therapeutic efficacy and dosimetry, as the progenies may accumulate in off-target tissues or regions. When comparing the 213Po peak in the cell sample and the supernatant, a shift to lower energies is observed in the cell sample spectrum. This shift is attributed to the presence of cells in the sample, which cause additional energy loss for alpha particles, whereas such interactions are absent in the supernatant. The reference sample, which contains no cells, exhibits very low activity, suggesting that the washing procedure effectively removed free activity that was not taken up by the cells.
@@
Figure 8: Alpha spectra of [225Ac]Ac-crown-TATE obtained from the cell sample (black), supernatant (blue), and reference sample (red, no cells), and (a): Linear scale. (b): Logarithmic scale.
@@ 3.4 Alpha spectrum from cell sample incubated with [225Ac]Ac3+
The uptake of free [225Ac]Ac3+ in AR42J cells was studied and the alpha spectra of cells alongside supernatant and the reference (no cells) were acquired using the BAD detection system as shown in 9(a). These plots indicate that there is uptake of free [225Ac]Ac3+ in the cells but less than [225Ac]Ac-crown-TATE. By comparing the spectra of the cell sample and the supernatant, it is evident that free [225Ac]Ac3+ is present in the supernatant, suggesting partial uptake of the isotope by the cells. As expected, the reference sample, which contains no cells, shows minimal activity, confirming that the washing procedure effectively removes unbound activity.
Figure 9: Alpha spectra of free [225Ac]Ac3+ obtained from the cell sample (black), supernatant (blue), and reference sample (red, no cells), and (a): Linear scale. (b): Logarithmic scale.
As a negative control, the behavior of [225Ac]Ac-PSMA-617 on AR42J cell cultures was investigated. 10(a) shows the alpha spectrum of [225Ac]Ac-PSMA-617 from the cell sample, along with the reference and cell supernatant samples, acquired using the BAD detection system in linear scale, while 10(b) presents the spectrum on a logarithmic scale. Since AR42J cells lack receptors for the PSMA-617 ligand, minimal uptake of [225Ac]Ac-PSMA-617 by the cells was expected. The alpha spectrum from the cell sample is almost identical to that of the supernatant, suggesting that [225Ac]Ac-PSMA-617 remained largely unbound and free in the medium, with minimal cellular uptake. The reference sample, which lacks cells, shows very low activity, confirming that the washing procedure effectively removed unbound [225Ac]Ac-PSMA-617, emphasizing the absence of significant binding to AR42J cells. The different shape of the 213Po peak in the supernatant and cell sample is related to the presence of cells, which causes an energy shift and broadening of the spectrum.
@@ 3.6 Geant4 simulation
Figure 11 shows a comparison between the experimental and simulated spectra for 225Ac distributed on Mylar foil near the detector. The four characteristic peaks of the ${}^{225}Ac$ decay chain ${}^{225}Ac$ (5830 keV), 221Fr (6340 keV), 217At (7067 keV), and 213Po (8376 keV) are clearly distinguishable. The simulation and experimental data are well-aligned, with a particularly strong match at the 213Po peak. The slight discrepancies observed in the other peaks may be attributed to minor differences in the physical setup, such as detector response or material variations in the sample cup that were not fully captured in the simulation.
@@
Figure 11: Comparison of experimental and simulated alpha spectra for dried 225Ac placed on Mylar foil near the detector.
@@ 4 Conclusion and Outlook
To the best of our knowledge, this study is the first to use an alpha detector for direct alpha spectroscopy of cancer cell samples under ambient conditions, demonstrating the potential of the Bio-Sample Alpha Detector (BAD) as a novel method in this context. The method enabled direct measurement of alpha spectra from AR42J rat pancreatic tumor cells incubated with [225Ac]Ac-crown-TATE and [225Ac]Ac-PSMA-617 and [225Ac]Ac3+. Results indicate significant uptake of [225Ac]Ac-crown-TATE by the cells, evidenced by distinct alpha spectra and minimal activity in the supernatant, suggesting efficient retention. In contrast, less efficient uptake of 225Ac-PSMA-617 was observed, correlating with the lack of PSMA-617 receptors on AR42J cells.
@@
This work underscores BAD’s versatility in conducting sensitive, real-time measurements of radiopharmaceuticals in cancer cells, enhancing understanding of radionuclide behavior in biological environments.
Future studies can explore additional cell types, radiolabeled compounds, and improved detector designs. Integration of a CO2 incubator and multi-channel data acquisition system will facilitate simultaneous measurements for statistically relevant systematic studies. A peristaltic pump system to control radiolabeled compound flow through cell cultures will simulate vascular conditions, providing insights into in vivo behavior of Targeted Alpha Therapy (TAT) agents. This study serves as an introduction to the BAD detector and method as an option for direct alpha spectroscopy. However, further optimization of the detector’s behavior for actual cell studies is essential to fully harness its potential for advancing TAT research and clinical applications.

### v2

@@ 2.5.3 Direct alpha spectroscopy of sell cultures incubated with labeled activity
Three sets of sample cups were prepared as described in Section 2.5. The first set contained [225Ac]Ac-crown-TATE, the second set contained [225Ac]Ac-PSMA-617 as a negative control (assuming the AR42J cells do not have receptors for PSMA-617), and the third set was incubated with [225Ac]Ac3+. For each set, a corresponding series of reference sample cups was prepared with the same amount of labeled activity and following the identical procedure, but without cells. Each sample cup was placed on the detector, and spectra were acquired for 900 seconds, achieving a statistical uncertainty of less than 1% in the total recorded counts. The experiment was repeated three times, and the standard deviation was applied to the results. Figure 2 illustrates a schematic view of the experimental setup.
@@
Figure 2: Schematic view of the experimental setup for direct alpha spectroscopy, showing the arrangement of AR42J cells with an approximate diameter of 13 µm. The cells are placed on a 2.5 µm Mylar foil, with a 100 µm air gap separating the cell sample and the Si-PIN diode (300 µm thickness). The media layer above the cells is approximately 3 mm thick. The yellow arrows represent the trajectory of emitted alpha particles.
@@ 2.5.4 Geant4 simulation
The silicon photodiode detector used in the simulation had dimensions of 18 $\times$ 18 mm and a thickness of 300 $\mu$m. A cylindrical sample cup with a diameter of 13 mm and a height of 3 mm was separated by a 100 $\mu$m air gap from the detector surface. A 2.5 $\mu$m thick Mylar foil was placed at the bottom of the sample cups. To model the radioactive source, 225Ac atoms were randomly distributed within a defined cylindrical volume (using the G4Tubs class), either as a thin layer on the Mylar foil or within a 13 $\mu$m-thick water layer representing a cell monolayer. The position of each 225Ac nucleus was assigned randomly within this volume using uniform random distributions (G4UniformRand() function), approximating a uniform source distribution. Only the dominant alpha emissions from 225Ac and its primary daughters were simulated; weaker decay branches were not considered. Each simulation was performed with $10^{6}$ primary events to ensure sufficient statistical accuracy. The simulated spectra for the Mylar surface and 13 $\mu$m water layer cases are presented in Figure 9 and Figure 10, respectively, alongside experimental data for comparison. The weaker decay branches were not considered in this simulation.
@@ 3.1 Alpha decay spectrum from 225Ac calibration source
Figure 3 shows the alpha spectrum for 225Ac obtained using the BAD detection system, revealing four distinct peaks for 225Ac (5830 keV), 221Fr (6340 keV), 217At (7067 keV), and 213Po (8376 keV). Calibration was performed using the high-energy flank of each peak. The characteristic tailing toward lower energies is attributed to alpha particle energy loss as they pass through both the 2.5 $\mu$m Mylar foil and the $\sim$100 $\mu$m air gap between the sample and the detector surface. The air gap dimension was determined from design drawings, as it could not be directly measured after assembly. The energy loss in air (12 keV per 100 $\mu$m) is relatively minor compared to the larger energy loss in the Mylar foil (260 keV for 2.5 $\mu$m thickness). Importantly, because the detector was calibrated using a dried 225Ac source placed on the Mylar within the final sample cup geometry, any energy loss contributions from both the Mylar and air gap are already incorporated into the calibration. Thus, small deviations between the design and physical assembly would not significantly affect the measurement accuracy. The overall detector efficiency across the entire energy range was calculated to be 37%.
@@ 3.2 Cell culture distribution
Figure 4 shows the distribution of AR42J cells on the Mylar foil. The cells appear to be evenly distributed across the surface of the sample cup, with no visible clumping or aggregation. This even distribution is crucial for ensuring uniform exposure during alpha spectroscopy. Microspheres with a known diameter of 20.1 $\mu$m were used as a reference, and the cell diameter was determined based on these microspheres using a microscope image and ImageJ software. A Gaussian fit was applied to the distribution, resulting in a cell diameter of 13.16 $\pm$ 0.12 $\mu$m. 5(a) shows the distribution of cell diameters in comparison with the microspheres. The size of the cells is important because it affects the ability of alpha particles to exit the cells and be detected before losing significant energy. An SRIM (Stopping and Range of Ions in Matter) simulation was performed to calculate the range of alpha particles in a water-based material, approximating the composition of the biological cells. The simulation included alpha particles emitted from 225Ac and its alpha-emitting daughters, using their characteristic energies. 5(b) shows the Bragg peak from the SRIM simulation, indicating that alpha particles from 225Ac have a range of approximately 50 $\mu$m in water. Considering that the average AR42J cell diameter is around 13 $\mu$m, the simulation confirms that alpha particles can escape the cells and reach the detector. This capability is essential for directly measuring the uptake and behavior of alpha-emitting radiopharmaceuticals.
@@
Refer to captionRefer to caption
Figure 5: (a) Normalized distribution of microspheres and AR42J cells based on diameter, with an inset showing a microscopic image of cells with reference spheres. Microspheres (black) have a diameter of 20 $\mu$m, while cells (red) show a wider distribution centered around 13.16 $\pm$ 0.12 $\mu$m. (b) Stopping power curves (Bragg peaks) simulated in SRIM for $\alpha$-particles emitted by 225Ac and its decay daughters in water. The 5.83 MeV $\alpha$-particles from 225Ac exhibit a range from approximately 40 $\mu$m to 50 $\mu$m, indicating that particles originating inside AR42J cells ($\sim$13 $\mu$m in diameter) can exit the cells and be detected.
@@ 3.3 Alpha spectrum from cell sample incubated with [225Ac]Ac-crown-TATE
6(a) presents the alpha spectrum of [225Ac]Ac-crown-TATE in the cell sample, alongside the corresponding reference and post-wash supernatant samples, acquired using the BAD detection system. The same spectra are shown on a logarithmic scale in 6(b) for better visualization of lower-intensity peaks. Although the detector resolution is insufficient to fully resolve closely spaced peaks such as those of 225Ac and 221Fr, peaks consistent with 217At and 213Po remain identifiable at their expected energy positions. The number of counts was normalized to the initial activity measured by a calibrated HPGe detector, enabling direct comparison between the three sample types. Three replicates of each sample were prepared, and the standard deviation is reflected in the plot. As observed in the spectrum from the cell sample, the alpha peaks appear broader compared to the dried sample, which is due to energy loss as alpha particles traverse through cellular material in aqueous media before reaching the detector. Interestingly, a slight energy shift of the 213Po peak is seen in the cell sample compared to the post-wash supernatant. While this shift might suggest increased energy loss in the presence of cells, the exact mechanism behind this observation remains uncertain, and no definitive conclusion can be drawn without further investigation. The supernatant shown in this figure refers to the media collected after the final wash step, following trypsinization and centrifugation, and thus contains minimal residual unbound activity. This signal corresponds to approximately $\sim$1.17of the cell-associated signal across the entire energy range. The 213Po peak is visible in the supernatant, likely reflecting free 213Bi released from the radiopharmaceutical complex due to the nuclear recoil of 225Ac decay. As previously described, such recoil can disrupt the coordination bonds of the labeled compound and cause daughter radionuclides to migrate into the surrounding solution. The reference sample, which contained no cells and underwent the same radiolabeling and washing procedures, exhibited only minimal detectable counts in the alpha spectrum, corresponding to approximately $\sim$0.09of the signal from the cell-associated fraction. This indicates that unbound radiolabeled compounds were effectively removed during sample preparation.
@@
Figure 6: Alpha spectra of [225Ac]Ac-crown-TATE obtained from the cell sample (black), supernatant (blue), and reference sample (red, no cells) (a): Linear scale. (b): Logarithmic scale.
As a negative control, the behavior of [225Ac]Ac-PSMA-617 on AR42J cell cultures was investigated. 7(a) shows the alpha spectrum of [225Ac]Ac-PSMA-617 from the cell sample, along with the reference and cell supernatant samples, acquired using the BAD detection system in linear scale, while 7(b) presents the spectrum on a logarithmic scale. Since AR42J cells lack receptors for the PSMA-617 ligand, minimal uptake of [225Ac]Ac-PSMA-617 by the cells was expected. The alpha spectrum from the cell sample is almost identical to that of the supernatant. Comparison of the net area under the spectra across the entire energy range indicates that the supernatant signal corresponds to approximately $\sim$79.5of the cell-associated signal, suggesting that[225Ac]Ac-PSMA-617 remained largely unbound and was released into the medium following centrifugation. The reference sample, which lacks cells, exhibits very low counts in the spectrum, approximately $\sim$0.55of the cell-associated signal, indicating that the washing procedure effectively removed unbound [225Ac]Ac-PSMA-617. Additionally, as mentioned in the previous section, the observed energy shift in the cell sample may be attributed to interactions with the cells and requires further investigation.
@@ 3.5 Alpha spectrum from cell sample incubated with [225Ac]Ac3+
The uptake of free [225Ac]Ac3+ in AR42J cells was studied, and the alpha spectra of the cell sample, supernatant, and reference (no cells) were acquired using the BAD detection system, as shown in 8(a) and 8(b). These spectra indicate that free [225Ac]Ac3+ is taken up by the cells, but to a lesser extent than [225Ac]Ac-crown-TATE (approximately 76% of the total cell-associated signal). By comparing the spectra of the cell sample and the supernatant, it is evident that free [225Ac]Ac3+ remains in the supernatant, with the net area under the supernatant spectrum across the entire energy range accounting for 40.1% of the cell sample signal. This finding suggests partial uptake of the isotope by the cells. As expected, the reference sample, which contains no cells, shows minimal activity (approximately 0.25% of the cell sample signal) confirming that the washing procedure effectively removes unbound activity.
Figure 8: Alpha spectra of free [225Ac]Ac3+ obtained from the cell sample (black), supernatant (blue), and reference sample (red, no cells) (a): Linear scale. (b): Logarithmic scale.
@@ 3.6 Geant4 simulation
Figure 9 shows a comparison between the experimental and simulated spectra for 225Ac distributed on the Mylar foil near the detector. The four characteristic peaks of the ${}^{225}Ac$ decay chain ${}^{225}Ac$ (5830 keV), 221Fr (6340 keV), 217At (7067 keV), and Polonium-213 (213Po) (8376 keV) are clearly distinguishable. The simulation and experimental data are well-aligned, with a particularly strong match at the 213Po peak. The slight discrepancies observed in the other peaks may be attributed to minor differences in the physical setup, such as detector response or material variations in the sample cup that were not fully captured in the simulation.
@@
Figure 9: Comparison of experimental and simulated alpha spectra for dried 225Ac placed on the Mylar foil near the detector.
@@ 4 Conclusion and outlook
To the best of our knowledge, this study is the first to employ an alpha detector for direct alpha spectroscopy of cancer cell samples under ambient conditions, demonstrating the potential of the Bio-sample Alpha Detector (BAD) as a novel method in the context of targeted alpha therapy. Presented as a proof-of-principle, this study enabled the direct measurement of alpha spectra from AR42J rat pancreatic tumor cells incubated with [225Ac]Ac-crown-TATE, [225Ac]Ac-PSMA-617, and [225Ac]Ac3+. The results were interpreted qualitatively and indicate significant uptake of [225Ac]Ac-crown-TATE by the cells, as evidenced by distinct alpha spectral features and minimal counts in the supernatant, suggesting efficient retention. In contrast, limited uptake of [225Ac]Ac-PSMA-617 was observed, consistent with the absence of PSMA receptors on AR42J cells.
@@
This paper serves as an initial validation of the BAD system and introduces it as a potential tool for studying alpha-emitting radiopharmaceuticals in vitro. In future work, we plan to develop a multi-channel BAD array (e.g., a 12-channel detector), integrated with a CO2 incubator and automated flow system, to enable quantitative, high-throughput studies. These improvements will allow for higher statistical power and more systematic evaluations of uptake, retention, and redistribution under physiologically relevant conditions. A peristaltic pump system to control radiolabeled compound flow through cell cultures will simulate vascular conditions, providing insights into in vivo behavior of TAT agents. Ultimately, this approach aims to provide a robust analytical framework for advancing TAT research and improving clinical translation of alpha-emitting therapeutics.